# Dynamic Degradation Decomposition Network for All-in-One Image Restoration

### v1

@@ IV Experiments
In this section, to validate the effectiveness of our method, we first introduce the experimental settings, followed by the qualitative and quantitative comparison results with the state-of-the-art (SOTA) approaches. Lastly, we report the results of ablation studies.
@@
Figure 4: Qualitative comparison with the SOTA approaches on the Rain100L, and GoPro datasets (better viewed by zooming in).
Figure 5: Qualitative comparison with the SOTA approaches on the BSD68, SOTS, and LOL datasets (better viewed by zooming in).
@@ IV-A Experimental Settings
Experimental Details. We implement our model under the PyTorch framework and train it on four NVIDIA Geforce RTX 4090 GPUs. The model employs the Adam optimizer with exponential decay rates of $\beta_{1}=0.9$ and $\beta_{2}=0.999$, trained for 1200 epochs. The initial learning rate is set to $1e^{-4}$ and reduces to $1e^{-6}$ through a cosine annealing strategy. During training, $128\times 128$ patches are randomly cropped from the original images as network inputs.
@@ IV-C1 Ablation Study of D3Net Components
We conduct a series of ablation experiments, as shown in Table IV. We sequentially remove Fourier Transform Analysis (FA), Cross-Attention Interaction (CA), Degradation Correction Prompt (DCP), Strategy Prompt (SP), and Decision Units (DUs), assessing the influence of each component to the performance of D3Net. We report the average PSNR and SSIM values of different methods on five image degradation tasks. These results lead us to four key observations:
@@ V Conclusion
In this paper, we propose a novel all-in-one image restoration network, termed D3Net, designed as a comprehensive architecture capable of handling multiple image restoration tasks. Distinct from existing all-in-one image restoration methods, the uniqueness of D3Net lies in its ability to guide the network learning process by deeply interacting frequency domain degradation characteristics with spatial domain image features to generate degradation correction prompt and strategy prompt, endowing the model with degradation adaptability. Moreover, we introduce a prompt-based dynamic decomposition mechanism for adaptable degradation decomposition. Owing to its adaptive prompt-guided mode, this mechanism can achieve superior degradation decomposition effects while reducing computational overhead. Extensive experiments demonstrate the effectiveness and scalability of the proposed D3Net.
Considering the reliance of the dynamic decomposition mechanism on the quality of prompt, we will explore a more robust and accurate mechanism for generating prompt in the future, such as incorporating external knowledge to mitigate the dependence on supervised data.

### v2

@@ IV Experiments
In this section, we first introduce the experimental settings, followed by the qualitative and quantitative comparison results with the state-of-the-art (SOTA) approaches. Lastly, we report the results of ablation studies.
@@
Figure 4: Qualitative comparison with the SOTA methods on the Rain100L, and GoPro datasets (better viewed by zooming in).
Figure 5: Qualitative comparison with the SOTA methods on the BSD68, SOTS, and LOL datasets (better viewed by zooming in).
@@ IV-A Experimental Settings
Experimental Details. We implement our model using the PyTorch framework and train it on four NVIDIA Geforce RTX 4090 GPUs. The model employs the Adam optimizer with exponential decay rates of $\beta_{1}=0.9$ and $\beta_{2}=0.999$, trained for 1200 epochs. The initial learning rate is set to $1e^{-4}$ and reduces to $1e^{-6}$ through a cosine annealing strategy. During training, $128\times 128$ patches are randomly cropped from the original images as network inputs.
@@ IV-C1 Ablation Study of D3Net Components
We conduct a series of ablation experiments, as shown in Table IV. We sequentially remove Fourier Transform Analysis (FA), Cross-Attention Interaction (CA), Degradation Correction Prompt (DCP), Strategy Prompt (SP), and Decision Units (DUs), assessing the influence of each component on the performance of D3Net. We report the average PSNR and SSIM values of different methods on five image degradation tasks. These results lead us to four key observations:
@@ V Conclusion
In this paper, we propose a novel all-in-one image restoration network, termed D3Net, as a comprehensive architecture capable of handling multiple image restoration tasks. Distinct from existing all-in-one image restoration methods, the uniqueness of D3Net lies in its ability to guide the network learning process by deeply interacting frequency domain degradation characteristics with spatial domain image features to generate degradation correction prompt and strategy prompt, endowing the model with degradation adaptability. Moreover, we introduce a prompt-based dynamic decomposition mechanism for adaptable degradation decomposition. Owing to its adaptive prompt-guided mode, this mechanism can achieve superior degradation decomposition effects while reducing computational overhead. Extensive experiments demonstrate the effectiveness and scalability of the proposed D3Net.
Considering the reliance of the dynamic decomposition mechanism on the quality of prompt, we will explore a more robust and accurate mechanism for generating prompt in the future.